OpenStudy (anonymous):

Evaluate g – 16, for g = 25

OpenStudy (anonymous):

9

OpenStudy (anonymous):

Assumption # 1: variable G = variable g if G = g = 25 then g - 16 = 25 - 16 = 5^2 - 4^2 = 3^2 = 9

OpenStudy (anonymous):

can you show steps so I can understand how you got that answer

OpenStudy (anonymous):

g-25, g-16 is 25-16 = 9

OpenStudy (anonymous):

Did u get t?

OpenStudy (anonymous):

yes i did thank you